using L’Hˆopital’s in the following questions

1) I = limx-0 [1/x - 2/ln(1 + 2x)]
2) limx-0 [e^x + x]^1/x
3) limx-1+ (ln x)(tan(pi*x/2))
1
Expert's answer
2020-10-01T10:57:21-0400

"\lim\limits_{x\rightarrow0}(\frac{1}{x} -\frac{2}{ln(1+2x)})=\lim\limits_{x\rightarrow0}(\frac{ln(1+2x)-2x}{xln(1+2x)})=\lim\limits_{x\rightarrow0}\frac{0}{0}" "=\lim\limits_{x\rightarrow0}(\frac{(ln(1+2x)-2x)'}{(xln(1+2x))'})="

"\lim\limits_{x\rightarrow0}(\frac{(ln(1+2x)-2x)'}{(xln(1+2x))'})=\lim\limits_{x\rightarrow0}[\frac{\frac{2}{1+2x}-2}{ln(1+2x)+\frac{2x}{1+2x}}]=
\lim\limits_{x\rightarrow0}\frac{-4x}{(1+2x)ln(1+2x)+2x}="

"\lim\limits_{x\rightarrow0}\frac{0}{0}=" "\lim\limits_{x\rightarrow0}\frac{-(4x)'}{((1+2x)ln(1+2x)+2x)'}=" "\lim\limits_{x\rightarrow0}\frac{-4}{2ln(1+2x)+(1+2x)*\frac{2}{1+2x}+2}=\lim\limits_{x\rightarrow0}\frac{-4}{2ln(1+2x)+4}=-1"


"\lim\limits_{x\rightarrow0}(e^x+x)^{\frac{1}{x}}=\lim\limits_{x\rightarrow0}e^{ln(e^x+x)^\frac{1}{x}}" -we use equality a= elna

"\lim\limits_{x\rightarrow0}ln(e^x+x)^\frac{1}{x} = \lim\limits_{x\rightarrow0}\frac{ln(e^x+x)}{x} =" "\lim\limits_{x\rightarrow0}\frac{0}{0}"-we use equality log(ab) = blog(a)


"\lim\limits_{x\rightarrow0}\frac{ln(e^x+x)}{x} = \lim\limits_{x\rightarrow0}(ln(e^x+x)^\frac{1}{x} = \lim\limits_{x\rightarrow0}\frac{(ln(e^x+x))'}{x'} =" "\lim\limits_{x\rightarrow0}\frac{\frac{e^x+1}{e^x+x}}{1}=2" hence

"\lim\limits_{x\rightarrow0}(e^x+x)^{\frac{1}{x}}=e^2"


"\lim\limits_{x\rightarrow1+}lnx*tan(\frac{\pi*x}{2})=\lim\limits_{x\rightarrow1+}\frac{lnx}{cot(\frac{\pi*x}{2})}=
\lim\limits_{x\rightarrow1+}\frac{0}{0}" we use equality tan(x) = 1/cot(x)

"\lim\limits_{x\rightarrow1+}\frac{lnx}{cot(\frac{\pi*x}{2})} =\lim\limits_{x\rightarrow1+}\frac{(lnx)'}{(cot(\frac{\pi*x}{2}))'}=\lim\limits_{x\rightarrow1+}\frac{\frac{1}{x}}{-\frac{\pi}{2}*csc^2(\frac{\pi*x}{2})}=" "-\frac{2}{\pi}"


Answer:

"\lim\limits_{x\rightarrow0}(\frac{1}{x} -\frac{2}{ln(1+2x)})=-1"

"\lim\limits_{x\rightarrow0}(e^x+x)^{\frac{1}{x}}=e^2"

"\lim\limits_{x\rightarrow1+}lnx*tan(\frac{\pi*x}{2})=-\frac{2}{\pi}"